Write a balanced equation for the following reaction description.

Calcium hydroxide and hydrogen chloride react to form calcium chloride and water.

Please explain
Answer #1

description.

The given reaction is between calcium hydroxide - Ca(OH)2 and hydrogen chloride-HCl react to form calcium chloride -CaCl2 and water-H2O.

lets assume reaction of 1 molecule of Ca(OH)2 .

In one Ca(OH)2 molecule, 1 calcium atom is present.

calcium is present only in CaCl2 in the product side which contain 1 calcium.

Therefore, number of CaCl2 = 1

In one CaCl2 molecule, 2 chlorine atoms are present.

chlorine is present only in HCl in the reactant side which contain 1 chlorine.

Therefore number of HCl = 2*1 = 2

In one Ca(OH)2 molecule, 2 oxygen atoms is present.

oxygen is present only in H2O in the product side which contain 1 calcium.

Therefore, number of H2O= 2*1 =2

Hence, balanced equation for the reaction: 1 Ca(OH)2 + 2 HCl \rightarrow 1 CaCl2 + 2 H2O.